Give an example of an infinite set with finite number of limit points, with proper

justification.


1
Expert's answer
2021-06-11T02:46:13-0400

The set


"\Big\{\frac{1}{n}\big| n\in \mathbb{Z}^+\Big\} \cup\Big\{1+\frac{1}{n}\big| n\in \mathbb{Z}^+\Big\}"

is an infinite set, which has a finite limit.

We can see this directly or we can use the assertion of finding limits in calculus.


For:


"\Big\{\frac{1}{n}\big| n\in \mathbb{Z}^+\Big\}\\
\lim_{n \rightarrow \infty}{\frac{1}{n}} = 0"

While for:


"\Big\{1+\frac{1}{n}\big| n\in \mathbb{Z}^+\Big\}\\
\lim_{n \rightarrow \infty}{1+\frac{1}{n}} = 1"

Thus:


"\Big\{\frac{1}{n}\big| n\in \mathbb{Z}^+\Big\} \cup\Big\{1+\frac{1}{n}\big| n\in \mathbb{Z}^+\Big\}= \{0,1\}"

Which is finite.